OpenStudy (anonymous):

what is the slope of the line passing through the points (2,7) and (-1,3)? a.1/7 b.3/4 c.4/3 d.1/3

OpenStudy (amtran_bus):

Use the point slope formula, y-y1=m(x-x1) The M is your slope. You will have to use y2-y1/x2-x1 to get slope and plug it in for M.

OpenStudy (amtran_bus):

Oh, I see it just wants slope. Just use the last equation I gave :)

OpenStudy (amtran_bus):

Find your y2 value (which is 3) and subtract it from your Y1 value (which is 7) Divide that by your x2 value - X1 value. I will let you find x2 and x1.

OpenStudy (ericdl):

slope = (y2 - y1) / (x2 - x1) slope = (3-7) / (-1-2) slope = -4 / -3 slope = 4/3 FINAL ANSWER

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what is this one

OpenStudy (ericdl):

(-7,-3) (1,-4) Slope = y2-y1/x2-x1 Slope = (-4+3)/(1+7) Slope = -1/8 FINAL ANSWER

OpenStudy (anonymous):

what is this one

OpenStudy (anonymous):

what is this one b

OpenStudy (ericdl):

(0,1) (4,-2) slope = y2-y1/x2-x1 slope = -2-1/4-0 slope = -3/4 final answer

OpenStudy (ericdl):

(0,-1) (4,-4) slope = -4+1 / 4-0 slope = -3/4 final answer

OpenStudy (anonymous):

which graph is it